You live 25 kilometers from your friend's home. Your parents need to drive you, so how long does it take to get them ready? Use algebra in solving such problem?


1
Expert's answer
2021-11-19T00:54:41-0500

We have the parameters;

Distance between both houses, d=25km

Let time be "t(hr)" = "t"

Let speed be represented by "s(km\/hr)"="s"


The question asked how long it takes to get them ready? And if possible to solve it algebraically!

Yes, there's basically no life situations we cannot express in simple algebraic terms!

Using the formula

"Speed=\frac{distance}{time}"

As an expression to show how long it'll take to have my parents drive me to my friend's house which is "25km" away

"S(km/h)=\frac{d(km)}{t(h)}"

Where "d(km)=25km"

"s(km/h)=\frac{25km}{t(h)}"

Making "t" subject of the equation

"t(h)=\frac{25km}{s(km/h)}"

"=\frac{25}{s}hr"

Where "s" is the "speed".


Also, if we seek to know how long it takes for me to get my parents all set(ready) to carry me to my friends house which is 25km away from my house, it still boils down to how fast I get to help them do some stuffs that might make them lag in preparation.

Since speed is inversely proportional to time

which implies the more the "speed" to get them ready, the less the "time" it will take to get them ready and vice versa!

i.e "Speed ∝\frac{1}{time}"